OpenStudy (anonymous):

lim x aprroachign 1 ( x-1)/(sqrt(x) - 1) = -2?

OpenStudy (anonymous):

yo this is so hard omg

OpenStudy (abb0t):

LOOOOOL!!!

OpenStudy (loser66):

lol

OpenStudy (luigi0210):

As soon as you start being nice, someone will be happy to help you

OpenStudy (anonymous):

Well, if we substituted x for 0 we'd get 0/0, which is undeterminate. Therefore, you need to factor the function to see if there's any other way to get a solution. So, since you have a square root, I'd try to rationalize the denominator. We'd get: (x-1)/(sqrt(x) - 1) all of that times * (sqrt(x) + 1)/(sqrt(x) + 1). After multiplying it out, we'd get: (x-1)*(sqrt(x) + 1) all of that divided / (x - 1) The x-1s cancel out and you get that the limit as x approaches 1 of that function is equal to the (sqrt(x) + 1). Substituting the x for the 1, you get 2. And that's the answer.

OpenStudy (luigi0210):

@Matt24 No giving away answers and don't do their work for them

OpenStudy (anonymous):

Why? Sorry I'm new here. I thought that was the purpose(?)

OpenStudy (luigi0210):

No, we guide the asker.

OpenStudy (abb0t):

@Luigi0210 is right. They are in Calculus now and they should be able to solve the problem with guidance. Only after they truly show efforts and still do not understand should you proceed to give them the answer so they truly understand.

OpenStudy (anonymous):

Oh, damn. I thought doing it myself would help the person in trouble. At least I think I wouldn't mind getting an answer myself after asking, because usually going online is my last resort after having thought about the problem for a long time

OpenStudy (kinggeorge):

Just to put my two cents in, we do encourage everyone to help the asker along instead of just providing them with the solution (regardless of how much work you show). Unfortunately, most people are not like you @Matt24 in that they think about the problem first. Many users just post questions in the hopes of getting a free answer. Those are the user you should report, and ignore. Welcome to OpenStudy!

OpenStudy (anonymous):

Oh well, thank you KingGeorge! I'm sorry for messing up :/ By the way, is there a section for Economics or Social Sciences here? Or is there only maths?

OpenStudy (luigi0210):

@Loser66 I've noticed buddy :3

OpenStudy (anonymous):

I agree with Loser66 that probably trying to explain is more time-consuming than just giving the answer. What happens if I spend lots of time online to explain sth to sbdy and then he or she doesn't really care?

OpenStudy (luigi0210):

@Matt24 Just click at the top where it says "Find More Subjects"

OpenStudy (anonymous):

@abb0t @Luigi0210 Thank you! I'm going to check it out

OpenStudy (anonymous):

So it's not wrong if I give an answer?

OpenStudy (luigi0210):

@Matt24 it is.. but chances are if you do you will get reported and get in trouble by the mods and such.

OpenStudy (luigi0210):

If you mean just giving them out directly

OpenStudy (anonymous):

Oh lol it's a difficult system

OpenStudy (kinggeorge):

It happens that you spend a lot of time carefully explaining something, and don't get anything out of it, you just learn to forget it and move on. Also, if you're halfway through your explanation, trying to get an asker involved, and they refuse to do so, don't be afraid to cut your losses and simply ignore the question until they start participating.

OpenStudy (anonymous):

Oh ok. Thank you!

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1}\frac{ x-1 }{ \sqrt{x}-1 }\] it can be solved two ways 1. \[\lim_{x \rightarrow 1}\frac{ \left( \sqrt{x} \right)^{2} -1^{2}}{ \sqrt{x}-1 }\] \[\lim_{x \rightarrow 1}\frac{ \left( \sqrt{x}+1 \right)\left( \sqrt{x}-1 \right) }{ \sqrt{x}-1 }\] cancel the common and plug x=1 to get the solution.

OpenStudy (anonymous):

2. multiply the numerator and denominator by \[\sqrt{x}+1\] \[\lim_{x \rightarrow 1}\frac{ x-1 }{ \sqrt{x}-1 }*\frac{ \sqrt{x}+1}{ \sqrt{x} +1}\] solve and get the answer